# Dissipation-dispersion analysis of fully-discrete implicit discontinuous Galerkin methods and application to stiff hyperbolic problems

### v2

@@ 3.1.1 The case of DIRK-DG approximations
$w_{0}=1+\frac{h^{2}\mathrm{k}^{2}}{8}+\frac{ih^{3}\mathrm{k}^{3}}{72}+O(h^{4}),\quad w_{1}=-\frac{h^{2}\mathrm{k}^{2}}{6}-\frac{ih^{3}\mathrm{k}^{3}}{72}+O(h^{4}).$
@@ 3.2 Numerical investigation of dissipation and dispersion in DIRK-DG schemes
In the following, and throughout the paper, we refer to the Courant number as the quantity
$C=\max_{x}\left\lVert\mathbf{f}^{\prime}(\mathbf{u})\right\rVert\frac{\Delta t}{h},$ (3.26)
where $\left\lVert\mathbf{f}^{\prime}(\cdot)\right\rVert$ denotes the spectral radius of the Jacobian of the flux function. In particular, in explicit DG schemes the CFL stability condition imposes the choice of a time-step $\Delta t_{\text{CFL}}$ such that $C\leq 1/(2p+1)$. Additionally, we define
@@
which represents the ratio between the time-step $\Delta t_{\text{imp}}$ used by an implicit time-integration and the time-step required for explicit stability. When $r=1$, then explicit and implicit schemes are run with the same time-step.
@@
Table 1: $2$-stage DIRK method. It is second-order accurate for any choice of $\gamma$.
@@
[…]
@@
Table 3: $4$-stage third-order DIRK method.
@@
(a) $2$-stage second-order DIRK method of Table 3 with $\gamma=0.2$.
[…]
@@
Figure 2: We show the development of instability when the DIRK parameters do not yield A-stable schemes. Top row: The white area represents the stability regions of the DIRK methods of order $p+1=2,3$, whereas the $\star$ markers are obtained from the eigenvalues $\Omega_{l}$, $l=0,\dots,p$, of the space $\mathbb{P}_{p}$ DG approximation as function of $K$. Bottom row: Measure of dissipation (3.22) computed for several Courant numbers.
@@
We observe that, for the $2$-stage DIRK in Table 3, any choice of $\gamma$ satisfies second-order conditions. In particular, we consider $\gamma\geq 0.25$. As already observed, $\gamma=1-\sqrt{2}/2\approx 0.29289$ gives the L-stable DIRK method of Pareschi and Russo [41]. Values of $\gamma<0.25$ do not guarantee A-stability of the method, as showed in Figure 2(a). In the top panel, we plot the stability region of the $2$-stage DIRK in Table 3 with $\gamma=0.2$. The white area represents the stability region of the DIRK method which clearly does not contain the whole half-plane $\mathfrak{Re}(z)<0$, $z\in\mathbb{C}$. The $\star$ markers are the points $\mathcal{P}_{l}(K)=(\mathfrak{Re}(\Omega_{l}(K)\Delta t),\mathfrak{Im}(\Omega_{l}(K)\Delta t)$, $l=0,1$, in the complex plane, where $\Omega_{l}(K)$ is the $l$-th eigenvalue of the space $\mathbb{P}_{1}$ DG approximation obtained with the wave number $K=\mathrm{k}h$. Here, $\Delta t_{\text{imp}}$ is computed with $r=4.5$, and we observe that for some $K$ the points $\mathcal{P}_{l}(K)$ lie outside the stability region of the DIRK method. In the bottom panel of Figure 2(a), we show the measure of dissipation (3.22) for $r=3,4.5,9$. Notice that for small $r$ it is still possible to guarantee the stability of the scheme since the dissipation is negative.
[…]
@@
[…]
@@
Figure 3: We explore the sensitivity of second-order schemes to the choice of the parameter $\gamma$. Top row: Measure of dissipation (3.22) and dispersion (3.21) for second-order DIRK-DG schemes as function of $K$ and $\gamma$. Blue $\times$ markers and red $+$ markers denote the values of the DIRK parameter for which the measure is minimum and maximum, respectively. Bottom row: Dissipation and dispersion measures obtained with DIRK parameters $\gamma=0.25,1-\sqrt{2}/2,0.5$.
@@
[…]
@@
We compare all these schemes on the solution of the linear advection equation (3.1) with speed $a=1$ and initial conditions (3.25a), (3.25b) and (3.25c). See Figure 5. Here, $r=15$ both for second- and third-order schemes, so that we can compare them when they have the same ratio between the time-step used and the time-step for the explicit stability. The solutions are reproduced with $N=400$ cells. We observe that the numerical investigation of the diffusion and dispersion errors agrees with the properties of the numerical solutions we show. In fact, in Figure 5(a) we notice that $\gamma=0.5$ provides a very dispersive scheme, whereas $\gamma=1-\sqrt{2}/2$ slightly dampens the extrema of the solution. We also observe that the two third-order schemes, see Figure 5(b), provide almost equivalent approximations in terms of dissipation. However, since they are less dispersive than second-order schemes, they allow to estimate better the speed of propagation and to cure spurious oscillations around discontinuities. In conclusion, these results show that the main gain in considering $p=2$ lies in its much lower dispersion compared to $p=1$.
@@ 4.1 Freezing the nonlinearity with a low-order approximation of the space limiters
The terms $\Delta u_{j}^{\min}$ and $\Delta_{\min}u_{j}$ are also computed on the predictor solution $\mathbf{u}^{\star}$. In this way, the complete scheme would be linear with respect to the space limiters, and nonlinear only through the flux function. With this definition, $\phi_{j}^{\star}\in\{0,1\}$ and therefore it allows us to detect whether a cell is troubling or not.
@@ 4.1.1 The space-time first-order implicit predictor
where $\epsilon$ is a small quantity. Direct substitution into classical Newton’s iteration (4.12) leads to the residual problem
@@ 4.1.2 The space-time high-order semi-linearized implicit scheme
which is the space approximation at the $i$-th DIRK stage. We use the subscript $\text{lim}^{\star}$ to stress that this is computed by freezing the limiters on the predictor solution. This term is now fully linear with respect to the unknowns at time $t^{(i)}=t_{n}+c_{i}\Delta t$, unless the non linearity of the flux function $\mathbf{f}$. Therefore, we approximate the nonlinear system (4.7) with
@@ 5 Numerical simulations
Finally, we recall the definitions of the Courant number, $C$, and the ratio $r$ between the time-step used and the time-step required for explicit stability, as defined in equations (3.26) and (3.27). These parameters will be extensively used in the following sections to describe the regimes under which the schemes operate.
@@ 5.1.1 Convergence test
We test the numerical convergence rate of DIRK-DG schemes on the linear scalar conservation law (3.1) with speed $a=1$ on $[-1,1]$ with periodic boundary conditions, up to the final time $t=2$. The initial condition
@@
$h$ $N$ $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ $20$ $22.30e{-}$ 2 $2.51$ $22.82e{-}$ 2 $2.19$ $0.18$ $1.50$ $0.5$ $1.52$ $1.25$ $1.15$
$5.00\cdot 10^{-2}$ $40$ $34.25e{-}$ 3 $2.43$ $36.65e{-}$ 3 $2.09$ $24.73e{-}$ 2 $1.91$ $0.14$ $1.83$ $0.47$ $1.43$
$2.50\cdot 10^{-2}$ $80$ $48.78e{-}$ 4 $2.27$ $31.62e{-}$ 3 $2.04$ $21.28e{-}$ 2 $1.89$ $23.62e{-}$ 2 $1.95$ $0.14$ $1.77$
$1.25\cdot 10^{-2}$ $160$ $41.88e{-}$ 4 $2.22$ $44.11e{-}$ 4 $1.98$ $33.19e{-}$ 3 $2.00$ $39.10e{-}$ 3 $1.99$ $23.59e{-}$ 2 $1.93$
$6.25\cdot 10^{-3}$ $320$ $54.06e{-}$ 5 $2.21$ $59.68e{-}$ 5 $2.09$ $48.11e{-}$ 4 $1.97$ $32.27e{-}$ 3 $2.00$ $39.06e{-}$ 3 $1.99$
@@
$h$ $N$ $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ $20$ $22.35e{-}$ 2 $2.52$ $24.63e{-}$ 2 $2.00$ $0.3$ $1.21$ $0.67$ $1.19$ $1.5$ $0.61$
$5.00\cdot 10^{-2}$ $40$ $34.51e{-}$ 3 $2.38$ $21.16e{-}$ 2 $2.00$ $28.78e{-}$ 2 $1.76$ $0.24$ $1.49$ $0.67$ $1.16$
$2.50\cdot 10^{-2}$ $80$ $49.44e{-}$ 4 $2.26$ $32.88e{-}$ 3 $2.01$ $22.45e{-}$ 2 $1.84$ $26.79e{-}$ 2 $1.82$ $0.24$ $1.49$
$1.25\cdot 10^{-2}$ $160$ $42.05e{-}$ 4 $2.20$ $47.28e{-}$ 4 $1.98$ $36.17e{-}$ 3 $1.99$ $21.75e{-}$ 2 $1.96$ $26.78e{-}$ 2 $1.82$
$6.25\cdot 10^{-3}$ $320$ $54.55e{-}$ 5 $2.17$ $41.83e{-}$ 4 $1.99$ $31.57e{-}$ 3 $1.97$ $34.40e{-}$ 3 $1.99$ $21.75e{-}$ 2 $1.96$
@@
$h$ $N$ $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ $20$ $22.59e{-}$ 2 $2.47$ $28.87e{-}$ 2 $1.84$ $0.57$ $0.97$ $1.12$ $1.05$ $2.22$ $0.12$
$5.00\cdot 10^{-2}$ $40$ $35.06e{-}$ 3 $2.35$ $22.31e{-}$ 2 $1.94$ $0.18$ $1.66$ $0.45$ $1.30$ $1.12$ $0.99$
$2.50\cdot 10^{-2}$ $80$ $31.10e{-}$ 3 $2.19$ $35.80e{-}$ 3 $1.99$ $25.02e{-}$ 2 $1.84$ $0.14$ $1.73$ $0.44$ $1.34$
$1.25\cdot 10^{-2}$ $160$ $42.63e{-}$ 4 $2.07$ $31.45e{-}$ 3 $2.00$ $21.27e{-}$ 2 $1.98$ $23.58e{-}$ 2 $1.93$ $0.14$ $1.71$
$6.25\cdot 10^{-3}$ $320$ $55.74e{-}$ 5 $2.20$ $43.65e{-}$ 4 $2.00$ $33.24e{-}$ 3 $1.97$ $39.04e{-}$ 3 $1.98$ $23.57e{-}$ 2 $1.93$
[…]
@@
$h$ $N$ $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
$0.2$ $10$ $29.24e{-}$ 2 $0.15$ $0.53$ $1.35$ $2.32$
$0.1$ $20$ $21.35e{-}$ 2 $2.78$ $22.09e{-}$ 2 $2.81$ $0.2$ $1.38$ $0.48$ $1.51$ $1.21$ $0.94$
$5.00\cdot 10^{-2}$ $40$ $32.04e{-}$ 3 $2.73$ $33.69e{-}$ 3 $2.50$ $23.69e{-}$ 2 $2.45$ $0.13$ $1.89$ $0.46$ $1.38$
$2.50\cdot 10^{-2}$ $80$ $42.79e{-}$ 4 $2.87$ $45.08e{-}$ 4 $2.86$ $36.71e{-}$ 3 $2.46$ $22.67e{-}$ 2 $2.26$ $0.13$ $1.85$
$1.25\cdot 10^{-2}$ $160$ $53.91e{-}$ 5 $2.83$ $41.25e{-}$ 4 $2.02$ $48.84e{-}$ 4 $2.92$ $33.76e{-}$ 3 $2.83$ $22.59e{-}$ 2 $2.31$
$6.25\cdot 10^{-3}$ $320$ $66.18e{-}$ 6 $2.66$ $55.43e{-}$ 5 $1.21$ $41.08e{-}$ 4 $3.03$ $44.92e{-}$ 4 $2.93$ $33.80e{-}$ 3 $2.77$
$3.13\cdot 10^{-3}$ $640$ $62.75e{-}$ 6 $4.30$
$1.56\cdot 10^{-3}$ $1{,}280$ $74.33e{-}$ 7 $2.67$
@@
$h$ $N$ $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate

### v3

@@ 3.1.1 The case of DIRK-DG approximations
$\nu_{0}=1+\frac{h^{2}\mathrm{k}^{2}}{8}+\frac{ih^{3}\mathrm{k}^{3}}{72}+O(h^{4}),\quad\nu_{1}=-\frac{h^{2}\mathrm{k}^{2}}{6}-\frac{ih^{3}\mathrm{k}^{3}}{72}+O(h^{4}).$
@@
Remark 3.3.
We can estimate the order of convergence of the scheme by using the error components defined in (3.17) and the Taylor expansions mentioned above. Using the upper bound
$\|\mathcal{E}^{n}_{m}\|^{2}\lesssim\sum_{l=0}^{p}|w_{l}|^{2}\,|\lambda_{l}^{n}-e^{-i\mathrm{k}at^{n}}|^{2},$
we can estimate the order of convergence for each error term, $m=0,\ldots,p$. On a single time step, for $n=1$ and from the expansions $\lambda_{0}=1-i\mathrm{k}\Delta t-\frac{\mathrm{k}^{2}}{2}\Delta t^{2}+O(\Delta t^{3})$ and $\nu_{0}=1+O(h^{2})$, we get
$\lambda_{0}-e^{-i\mathrm{k}a\Delta t}=O(\Delta t^{3})\quad\mbox{and}\quad|\nu_{0}(\lambda_{0}-e^{-i\mathrm{k}a\Delta t})|=O(\Delta t^{3}).$
From $\lambda_{1}=O\Big(\frac{\Delta t}{h}\Big)+O\Big(\frac{\Delta t^{2}}{h^{2}}\Big)+O\Big(\frac{\Delta t^{3}}{h^{3}}\Big)+\dots$ and $\nu_{1}=O(h^{2})$, one obtains
$|\nu_{1}(\lambda_{1}-e^{-i\mathrm{k}a\Delta t})|\sim O(h^{2}).$
Therefore, with respect to the norm $\|\mathcal{E}^{n}_{m}\|$, $m=0,\ldots,p$ the dominant contributions are
$\|\mathcal{E}_{m}\|\sim O(h^{2})+O(\Delta t^{3}).$
Consequently, the method is of second order in space and third order in time. Furthermore: if $\Delta t\sim h$, the spatial error dominates, yielding overall order $2$; if $\Delta t\sim h^{2/3}$ so that the terms are balanced, the temporal order $\sim 3$ is achieved.
@@ 3.2 Numerical investigation of dissipation and dispersion in DIRK-DG schemes
Throughout the paper, we define
$\Delta t_{\mathrm{CFL}}:=\frac{h}{(2p+1)\,\max_{x}\left\lVert\mathbf{f}^{\prime}(\mathbf{u})\right\rVert},$ (3.26)
where $\left\lVert\mathbf{f}^{\prime}(\mathbf{u})\right\rVert$ denotes the spectral radius of the Jacobian of the flux function. This corresponds to the classical CFL restriction for a discontinuous Galerkin scheme of degree $p$ with a Rusanov flux. In general, the precise stability bound of an explicit scheme depends also on the particular Runge-Kutta integrator employed. However, in this work we take (3.26) as a reference definition of $\Delta t_{\mathrm{CFL}}$, independently of the specific Runge-Kutta scheme, in order to provide a consistent baseline for comparing explicit and implicit methods. We then introduce
@@
that is, the ratio between the implicit time-step $\Delta t_{\text{imp}}$ and the reference explicit stability limit $\Delta t_{\text{CFL}}$ running the same test with an explicit RKDG of matching order in space and time. When $r=1$, implicit and explicit schemes are advanced with the same time-step.
@@
Table 1: 22-stage DIRK method. It is second-order accurate for any choice of $\gamma$.
@@
Table 2: 33-stage DIRK method. It is third-order accurate for $\gamma=0.158983899988677$, $\gamma=0.435866521508459$, $\gamma=2.405149578502864$.
@@
Table 3: 44-stage third-order DIRK method.
@@
(a) 22-stage second-order DIRK method of Table 3 with $\gamma=0.2$.
(b) 33-stage third-order DIRK method of Table 3 with $\gamma=0.158983899988677$.
(c) 33-stage third-order DIRK method of Table 3 with $\gamma=2.405149578502864$.
@@
Figure 2: We show the development of instability when the DIRK parameters do not yield A-stable schemes. Top row: The white area represents the stability regions of the DIRK methods of order $p+1=2,3$, whereas the $\star$ markers are obtained from the eigenvalues $\Omega_{l}$, $l=0,\dots,p$, of the matrix coming from the space $\mathbb{P}_{p}$ DG approximation as function of $K$. Bottom row: Measure of dissipation (3.22) computed for several Courant numbers.
@@
We observe that, for the $2$-stage DIRK in Table 3, any choice of $\gamma$ satisfies second-order conditions. In particular, we consider $\gamma\geq 0.25$. As already observed, $\gamma=1-\sqrt{2}/2\approx 0.29289$ gives the L-stable DIRK method of Pareschi and Russo [41]. Values of $\gamma<0.25$ do not guarantee A-stability of the method, as showed in Figure 2(a). In the top panel, we plot the stability region of the $2$-stage DIRK in Table 3 with $\gamma=0.2$. The white area represents the stability region of the DIRK method which clearly does not contain the whole half-plane $\mathfrak{Re}(z)<0$, $z\in\mathbb{C}$. The $\star$ markers are the points $\mathcal{P}_{l}(K)=(\mathfrak{Re}(\Omega_{l}(K)\Delta t),\mathfrak{Im}(\Omega_{l}(K)\Delta t)$, $l=0,1$, in the complex plane, where $\Omega_{l}(K)$ depends on the $l$-th eigenvalue of the matrix coming from the space $\mathbb{P}_{1}$ DG approximation obtained with the wave number $K=\mathrm{k}h$. Here, $\Delta t_{\text{imp}}$ is computed with $r=4.5$, and we observe that for some $K$ the points $\mathcal{P}_{l}(K)$ lie outside the stability region of the DIRK method. In the bottom panel of Figure 2(a), we show the measure of dissipation (3.22) for $r=3,4.5,9$. Notice that for small $r$ it is still possible to guarantee the stability of the scheme since the dissipation is negative. We remark that the plots in the bottom panels of Figure 2 are restricted to $K\in[0,\pi]$ since the discrete Fourier spectrum is $2\pi$-periodic in $K$. Instead, in the top panels of Figure 2 we consider $K\in[0,2\pi]$ to show that the eigenvalue curves close, as typically expected.
For the $3$-stage DIRK scheme in Table 3, we observe that third-order conditions are satisfied for $\gamma=0.158983899988677$, $\gamma=0.435866521508459$ and $\gamma=2.405149578502864$. The second value allows to recover the A-stable and stiffly accurate, and therefore L-stable, method of [4]. We will consider only the value $\gamma=0.435866521508459$ since the other parameter values provide methods that are not A-stable. This is observed in the top panels of Figure 2(b) and Figure 2(c), where we plot the stability region of the $3$-stage DIRK in Table 3 with $\gamma=0.158983899988677$ and $\gamma=2.405149578502864$, respectively. As in the second-order DIRK, the stability region is represented by the white area which does not contain the whole half-plane $\mathfrak{Re}(z)<0$, $z\in\mathbb{C}$. The $\star$ markers are the points $\mathcal{P}_{l}(K)=(\mathfrak{Re}(\Omega_{l}(K)\Delta t),\mathfrak{Im}(\Omega_{l}(K)\Delta t)$, $l=0,1,2$, in the complex plane, where $\Omega_{l}(K)$ is the $l$-th eigenvalue of the matrix coming from the space $\mathbb{P}_{2}$ DG approximation obtained with the wave number $K=\mathrm{k}h$. Here, $\Delta t_{\text{imp}}$ is computed with $r=3$. The zoom shows that for some $K$ the points $\mathcal{P}_{l}(K)$ lie outside the stability regions of the DIRK methods. In the bottom panels of Figure 2(b) and Figure 2(c), we show the measure of dissipation (3.22) for $r=3,4.5,9$. Notice that the instability of the schemes are highlighted also by the negative dissipation.
The L-stable third-order method of [4], obtained with the parameter $\gamma=0.435866521508459$, is compared in the following with the $4$-stage third-order accurate method given in Table 3 which is a Strong Stability Preserving (SSP) scheme given in [48].
@@
${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ $\mathbb{P}_{2}$ DG coupled with the $3$-stage third-order DIRK in Table 3 with $\gamma=0.435866521508459$
@@
Figure 3: We explore the sensitivity of second-order schemes to the choice of the parameter $\gamma$. Top row: Measure of dissipation (3.22) and dispersion (3.21) for second-order DIRK-DG schemes as function of $K$ and $\gamma$. Blue $\times$ markers and red ++ markers denote the values of the DIRK parameter for which the measure is minimum and maximum, respectively. Bottom row: Dissipation and dispersion measures obtained with DIRK parameters $\gamma=0.25,1-\sqrt{2}/2,0.5$.
@@
In Figure 4, we present the dissipation and dispersion measures of two selected DIRK schemes of order 3. Specifically, we examine the L-stable DIRK scheme from Table 3 with $\gamma=0.435866521508459$ and the SSP DIRK scheme from Table 3. The results are displayed for three ratio numbers $r=15,25,50$. We observe that the $4$-stage SSP DIRK scheme of order 3 is slightly less diffusive and less dispersive than the $3$-stage L-stable DIRK scheme of order 3. Additionally, both dissipation and dispersion measures increase with increasing Courant number.
@@
We compare all these schemes on the solution of the linear advection equation (3.1) with speed $a=1$ and initial conditions (3.25a), (3.25b) and (3.25c). See Figure 5. Here, $r=15$ both for second- and third-order schemes, so that we can compare them when they have the same ratio between the time-step used and the time-step for the explicit stability. The solutions are computed with $N=400$ cells. We observe that the numerical investigation of the diffusion and dispersion errors agrees with the properties of the numerical solutions we show. In fact, in Figure 5(a) we notice that $\gamma=0.5$ provides a very dispersive scheme, whereas $\gamma=1-\sqrt{2}/2$ slightly dampens the extrema of the solution. We also observe that the two third-order schemes, see Figure 5(b), provide almost equivalent approximations in terms of dissipation. However, since they are less dispersive than second-order schemes, they allow to estimate better the speed of propagation and to cure spurious oscillations around discontinuities. In conclusion, these results show that the main gain in considering $p=2$ lies in its much lower dispersion compared to $p=1$.
@@ 4.1 Freezing the nonlinearity with a low-order approximation of the space limiters
The terms $\Delta u_{j}^{\min}$ and $\Delta_{\min}u_{j}$ are also computed on the predictor solution $\mathbf{u}^{\star}$. Specifically, $\Delta u_{j}^{\min}=u_{j}^{\star}-\min(u_{j-1}^{\star},u_{j}^{\star},u_{j+1}^{\star})$ and $\Delta_{\min}u_{j}=u_{j}^{\star}-\min(u_{h}^{\star}(x_{j+1/2},t),u_{h}^{\star}(x_{j-1/2},t))$. In this way, the complete scheme would be linear with respect to the space limiters, and nonlinear only through the flux function. With this definition, $\phi_{j}^{\star}\in\{0,1\}$ and therefore it allows us to detect whether a cell is troubling or not.
@@ 4.1.1 The space-time first-order implicit predictor
where $\epsilon$ is a small quantity, of order $\sim 10^{-5}$. Direct substitution into classical Newton’s iteration (4.12) leads to the residual problem
@@ 4.1.2 The space-time high-order semi-linearized implicit scheme
which is the space approximation at the $i$-th DIRK stage. We use the subscript $\text{lim}^{\star}$ to stress that this is computed by freezing the limiters on the predictor solution. This term is now fully linear with respect to the unknowns at time $t^{(i)}=t_{n}+c_{i}\Delta t$, unless the non-linearity of the flux function $\mathbf{f}$. Therefore, we approximate the nonlinear system (4.7) with
@@ 5 Numerical simulations
Finally, we recall the definition of the ratio $r$ between the time-step used and the time-step required for explicit stability, as defined in equation (3.27). This parameter will be extensively used in the following sections to describe the regimes under which the schemes operate.
@@ 5.1.1 Convergence test
We test the numerical convergence rate of DIRK-DG schemes on the linear scalar conservation law (3.1) with speed $a=1$ on [1,1][-1,1] with periodic boundary conditions, up to the final time $t=2$. The initial condition
@@
$h$ NN $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ 2020 2.30e22.30e{-}2 $2.51$ 2.82e22.82e{-}2 $2.19$ $0.18$ $1.50$ $0.5$ $1.52$ $1.25$ $1.15$
$5.00\cdot 10^{-2}$ $40$ 4.25e34.25e{-}3 $2.43$ 6.65e36.65e{-}3 $2.09$ 4.73e24.73e{-}2 $1.91$ $0.14$ $1.83$ $0.47$ $1.43$
$2.50\cdot 10^{-2}$ $80$ 8.78e48.78e{-}4 $2.27$ 1.62e31.62e{-}3 $2.04$ 1.28e21.28e{-}2 $1.89$ 3.62e23.62e{-}2 $1.95$ $0.14$ $1.77$
$1.25\cdot 10^{-2}$ $160$ 1.88e41.88e{-}4 $2.22$ 4.11e44.11e{-}4 $1.98$ 3.19e33.19e{-}3 $2.00$ 9.10e39.10e{-}3 $1.99$ 3.59e23.59e{-}2 $1.93$
$6.25\cdot 10^{-3}$ $320$ 4.06e54.06e{-}5 $2.21$ 9.68e59.68e{-}5 $2.09$ 8.11e48.11e{-}4 $1.97$ 2.27e32.27e{-}3 $2.00$ 9.06e39.06e{-}3 $1.99$
@@
$h$ NN $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ 2020 2.35e22.35e{-}2 $2.52$ 4.63e24.63e{-}2 $2.00$ $0.3$ $1.21$ $0.67$ $1.19$ $1.5$ $0.61$
$5.00\cdot 10^{-2}$ $40$ 4.51e34.51e{-}3 $2.38$ 1.16e21.16e{-}2 $2.00$ 8.78e28.78e{-}2 $1.76$ $0.24$ $1.49$ $0.67$ $1.16$
$2.50\cdot 10^{-2}$ $80$ 9.44e49.44e{-}4 $2.26$ 2.88e32.88e{-}3 $2.01$ 2.45e22.45e{-}2 $1.84$ 6.79e26.79e{-}2 $1.82$ $0.24$ $1.49$
$1.25\cdot 10^{-2}$ $160$ 2.05e42.05e{-}4 $2.20$ 7.28e47.28e{-}4 $1.98$ 6.17e36.17e{-}3 $1.99$ 1.75e21.75e{-}2 $1.96$ 6.78e26.78e{-}2 $1.82$
$6.25\cdot 10^{-3}$ $320$ 4.55e54.55e{-}5 $2.17$ 1.83e41.83e{-}4 $1.99$ 1.57e31.57e{-}3 $1.97$ 4.40e34.40e{-}3 $1.99$ 1.75e21.75e{-}2 $1.96$
@@
$h$ NN $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
@@
$0.1$ 2020 2.59e22.59e{-}2 $2.47$ 8.87e28.87e{-}2 $1.84$ $0.57$ $0.97$ $1.12$ $1.05$ $2.22$ $0.12$
$5.00\cdot 10^{-2}$ $40$ 5.06e35.06e{-}3 $2.35$ 2.31e22.31e{-}2 $1.94$ $0.18$ $1.66$ $0.45$ $1.30$ $1.12$ $0.99$
$2.50\cdot 10^{-2}$ $80$ 1.10e31.10e{-}3 $2.19$ 5.80e35.80e{-}3 $1.99$ 5.02e25.02e{-}2 $1.84$ $0.14$ $1.73$ $0.44$ $1.34$
$1.25\cdot 10^{-2}$ $160$ 2.63e42.63e{-}4 $2.07$ 1.45e31.45e{-}3 $2.00$ 1.27e21.27e{-}2 $1.98$ 3.58e23.58e{-}2 $1.93$ $0.14$ $1.71$
$6.25\cdot 10^{-3}$ $320$ 5.74e55.74e{-}5 $2.20$ 3.65e43.65e{-}4 $2.00$ 3.24e33.24e{-}3 $1.97$ 9.04e39.04e{-}3 $1.98$ 3.57e23.57e{-}2 $1.93$
Table 11: Experimental order of convergence of the third-order scheme ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$, computed on the initial condition (5.2) and the DIRK scheme in Table 3 with $\gamma=0.435866521508459$.
@@
$h$ NN $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate
$0.2$ $10$ 9.24e29.24e{-}2 $0.15$ $0.53$ $1.35$ $2.32$
$0.1$ 2020 1.35e21.35e{-}2 $2.78$ 2.09e22.09e{-}2 $2.81$ $0.2$ $1.38$ $0.48$ $1.51$ $1.21$ $0.94$
$5.00\cdot 10^{-2}$ $40$ 2.04e32.04e{-}3 $2.73$ 3.69e33.69e{-}3 $2.50$ 3.69e23.69e{-}2 $2.45$ $0.13$ $1.89$ $0.46$ $1.38$
$2.50\cdot 10^{-2}$ $80$ 2.79e42.79e{-}4 $2.87$ 5.08e45.08e{-}4 $2.86$ 6.71e36.71e{-}3 $2.46$ 2.67e22.67e{-}2 $2.26$ $0.13$ $1.85$
$1.25\cdot 10^{-2}$ $160$ 3.91e53.91e{-}5 $2.83$ 1.25e41.25e{-}4 $2.02$ 8.84e48.84e{-}4 $2.92$ 3.76e33.76e{-}3 $2.83$ 2.59e22.59e{-}2 $2.31$
$6.25\cdot 10^{-3}$ $320$ 6.18e66.18e{-}6 $2.66$ 5.43e55.43e{-}5 $1.21$ 1.08e41.08e{-}4 $3.03$ 4.92e44.92e{-}4 $2.93$ 3.80e33.80e{-}3 $2.77$
$3.13\cdot 10^{-3}$ $640$ 2.75e62.75e{-}6 $4.30$
$1.56\cdot 10^{-3}$ 1,2801{,}280 4.33e74.33e{-}7 $2.67$
@@
$h$ NN $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate $L^{1}$ error rate